# LASeR: Learning to Adaptively Select Reward Models with Multi-Armed Bandits

### v1

@@ Appendix B Additional empirical results
Table 4: Performance of 4 RMs including OLMo, Eurus, Zephyr and Qwen1.5.
Method Llama-3-8B Mistral-7B
StrategyQA GSM8K MMLU StrategyQA GSM8K MMLU
OLMo-7B-Instruct 80.23 68.91 65.74 68.73 44.96 56.94
Eurus-7b-kto 81.15 71.13 66.26 68.64 45.37 56.96
Zephyr-7b-alpha 84.29 73.16 67.15 70.06 45.81 62.04
Qwen1.5-7B-Chat 84.79 73.07 67.53 71.05 45.74 62.38
Method Llama-3-8B Mistral-7B
StrategyQA GSM8K MMLU Avg. StrategyQA GSM8K MMLU Avg.
SFT 80.41 69.43 65.66 71.83 68.57 43.62 56.48 56.22
Perplexity 80.55 69.21 65.62 71.79 68.83 43.47 57.14 56.48
Informativeness 82.87 73.55 66.69 74.37 70.29 44.98 59.29 58.19
Random RM Selection 82.72 70.93 66.10 73.25 69.24 44.05 57.68 56.99
Sequential RM Selection 83.15 73.38 66.17 74.23 70.40 44.79 59.07 58.09
LASeR (Ours) 83.54 73.80 66.79 74.71 70.91 44.93 59.63 58.49
Appendix C Generalization Capabilites of LASeR
LASeR Training with Domain-specific Evaluation Metrics. While recent works focus on building RMs that reflect preferences across domains, an extensive body of prior work develop a suite of evaluation metrics catered to specific domains such as reasoning (Golovneva et al., 2022; Prasad et al., 2023). To show that LASeR can be used to select any kind of evaluation metric from a collection of metrics during training, in Table 5, we present results with training LLMs on model-based metrics from ROSCOE (Golovneva et al., 2022) by replacing RMs with informativeness, faithfulness, reasoning alignment, hallucination, common sense error, semantic, coherence and perplexity in Sec. 3. Llama-3-8B models trained using LASeR yields 1.62% accuracy improvement over baselines on average across three datasets. These results are also generalized to Mistral-7B, except for GSM8K, where we achieve comparable performance to the Base + Informativeness baselines. Note that the perplexity of most responses is nearly identical, making it difficult to distinguish between them, explaining why perplexity shows little to no improvement compared to the base model.
Robustness to Underperforming Evaluation Metrics.
Similar to our analysis on noise in rewards in Sec. 5, we investigate how adding ROSCOE metrics with poor correlation to human-annotated labels in meta-evaluation by Golovneva et al. (2022) impacts the performance of Llama-3-8B on GSM8K. Once again, even with ROSCOE metrics, demonstrates LASeR can maintain consistent performance by adaptively prioritizing the most relevant reward signals, outperforming the sequential baseline, which fails to filter out irrelevant information effectively. Fig. 8 shows that as the number of irrelevant metrics increases, LASeR’s selection strategy continues to perform robustly. Specifically, LASeR has an average performance drop of only $0.13\%$, whereas the sequential baseline suffers a $2.15\%$ accuracy drop under the same conditions. Lastly, LASeR using Exp3 maintains a consistent performance level with a $0.4\%$ accuracy drop.
@@ Appendix C Generalization Capabilites of LASeR
Table 6: Across different training loss functions (Sec. 3.1), optimizing with multiple RMs via LASeR outperforms the sequential RM selection with Llama-3-8B.
Loss Method StrategyQA GSM8K MMLU Avg.
@@
LASeR 85.96 74.75 68.02 76.24
Generalization to Training Loss Functions. In Sec. 3.1, we state that the choice of loss function used to train the LLM depends on the underlying task or domain. Nevertheless, we always use the training loss as the MAB reward to update the MAB’s parameters. Here we study the performance of LASeR and baselines with 4 different loss functions, NLL, DPO, NLL + DPO (Pang et al., 2024), and KTO (Ethayarajh et al., 2024), in the reasoning domain. Results in Table 6 show that training LLMs with multiple rewards using LASeR outperforms sequential RM selection by $2.4\%$, $1.7\%$, and $1.3\%$ when using NLL, DPO, NLL+DPO loss functions, respectively; while both methods yield comparable performance with KTO. Additionally, we found that the most effective training loss functions are NLL + DPO for StrategyQA, NLL for GSM8K, and KTO for MMLU. However, irrespective of the choice of the underlying loss function, LASeR is more effective at balancing and adaptively selecting from multiple RMs.
@@
Method Llama-3-8B Mistral-7B
CSQA MATH Avg. CSQA MATH Avg.
SFT 65.64 29.13 47.39 59.06 16.38 37.72
Best RM 67.59 31.54 49.57 60.46 18.08 39.27
Avg. RM 67.16 30.36 48.76 60.06 17.25 38.66
Random RM Selection 68.31 30.21 49.26 60.19 16.96 38.58
Sequential RM Selection 67.73 30.25 48.99 60.56 17.96 39.26
LASeR (Ours) 69.26 31.67 50.47 61.65 18.97 40.31
Generalization to OOD Tasks.
We first assess the generalization ability of our method by training models on specific datasets and evaluating their performance on out-of-distribution reasoning tasks. Specifically, we train the model on the StrategyQA and MMLU datasets and evaluate its generalization on the CommonsenseQA (CSQA; Talmor et al., 2019) dataset. Similarly, we train on GSM8K and test on MATH (Hendrycks et al., 2021c) to assess the model’s ability to generalize across different reasoning datasets. These tasks are designed to capture both general reasoning ability and OOD generalization across domains. We report the results in Table 7, where we find that across both Llama-3-8B and Mistral-7B models, models trained with LASeR yield the best average accuracy beating training with the best RM by roughly $2\%$ (absolute) on CSQA with Llama-3-8B. On Mistral-7B, training with LASeR outperforms both training with single best RM and sequential RM selection by slightly over $1\%$.

### v2

@@ Appendix B Additional Empirical Results
Table 6: Performance of 4 RMs including OLMo, Eurus, Zephyr, and Qwen1.5 on Llama-3-8B. The best is bolded, and the second-best is underlined.
Method StrategyQA GSM8K MMLU Avg.
OLMo-7B-Instruct 80.23 68.91 65.74 71.62
Eurus-7B-Kto 81.15 71.13 66.26 72.84
Zephyr-7B-Alpha 84.29 73.16 67.15 74.87
Qwen1.5-7B-Chat 84.79 73.07 67.53 75.13
LASeR (Ours) 85.96 74.75 68.24 76.32
Table 7: Win rates of LASeR against individual RMs.
Win Rates
LASeR vs. Zephyr 56.34%
LASeR vs. Qwen 63.47%
LASeR vs. Eurus 59.02%
LASeR vs. Olmo 66.72%
Impact of RMs Diversity.
In this experiment, we test the impact of RMs diversity on downstream performance. Instead of selecting RMs based on overall RewardBench scores across categories as in our initial experiments, we created a more capability-diverse RM set by choosing the best-performing RM in each category in RewardBench: "Chat", "Chat Hard", "Safety", and "Reasoning". This selection improves the diversity in the types of tasks and evaluation criteria each RM specializes in. The new RM set includes Qwen1.5-7B-Chat, Zephyr-7B-Alpha, Tulu-2-DPO-7B, and MMPO-Gemma-7B, replacing Eurus and OLMo models from the previous RM set. Using this updated RM set, we test LASeR’s ability (denoted as "LASeR-Diverse") to generalize across diverse categories on WildChat. We compare LASeR-Diverse to the Best RM (from the revised RM set) and RM Agreement Ensemble baselines (the strongest baselines across our benchmarks) and report the results in Table 8. These results show that with a new set of RMs, LASeR continues to perform better, outperforming both the Best RM and RM Agreement Ensemble baselines. We emphasize that whether RMs are selected based on overall scores or diversity across RewardBench categories, LASeR consistently outperforms baselines. This reflects the core contribution of LASeR: its ability to automatically select reward models at the instance level, achieving consistent gains across tasks without requiring manual tuning or oracle selection. Users can simply provide a pool of RMs without knowing which is best.
Table 8: Win rates of LASeR against the two best baselines with a more diverse set of RMs.
Win Rates
LASeR-Diverse vs. Best RM 56.28%
LASeR-Diverse vs. RM Agree. Ensemble 57.21%
Appendix C Generalization Capabilities of LASeR
Generation to other Bandit Algorithms.
To demonstrate the generalization of LASeR to other Bandit Algorithms, we include empirical comparisons with NeuralUCB, a more expressive contextual bandit and Exp3, a widely used non-contextual bandit algorithm. NeuralUCB is parameterized by a 3-layer MLP with hidden dimensions 256 and 128. The output layer of NeuralUCB produces a single scalar, which is the predicted expected reward for each RM based on the context. Table 9 shows that LinUCB achieves comparable performance to Exp3 and performs slightly below NeuralUCB on 2 out of 3 datasets. However, LinUCB is significantly more efficient and stable because it uses closed-form updates and maintains simple per-arm information, avoiding the overhead of training neural networks as in NeuralUCB. These results support our central claim: regardless of the specific bandit algorithm used, LASeR consistently outperforms baseline RM selection and ensemble baselines.
Table 9: Performance and efficiency of LASeR with different bandit algorithms.
Method StrategyQA GSM8K MMLU Wall-clock hours (Avg.)
LASeR (LinUCB) 85.96 74.75 68.24 5.92
LASeR (Exp3) 85.58 75.04 68.56 6.26
LASeR (NeuralUCB) 85.73 75.62 69.15 10.27
Generalization to Different Base Models.
In the main paper, we present the results for the Llama-3-8B model. Here we present the results for Mistral-7B model across reasoning, instruction-following and long-context understanding tasks.
In reasoning tasks, LASeR improves average accuracy by roughly $2\%$ and $1.44\%$ over the sequential baseline and the best RM baseline, respectively. Compared to ensemble baselines, LASeR outperforms RM Agreement Ensemble, Online RM Ensemble, and RM Score Ensemble by $1.65\%$ and $2.71\%$, respectively. Overall, we observe the same results as Llama-3-8B model, we achieve consistent results while the underlined second-place models show inconsistent performance across datasets and models.
Table 10: Performance on reasoning benchmarks on Mistral-7b. The baselines also include supervised fine-tuning on human-written responses (SFT) as a reference for performance without preference optimization. The highest accuracy is shown in bold, and the second-highest accuracy is underlined. LASeR yields the highest average accuracy.
Method StrategyQA GSM8K MMLU Avg.
SFT 68.57 43.62 56.48 56.22
Best RM 70.06 45.81 62.04 59.30
Avg. RM 69.62 45.47 59.58 58.22
Random RM Selection 69.97 45.12 59.88 58.32
Seq. RM Selection 70.59 46.11 59.66 58.79
Classifier Selection 70.31 45.28 60.35 58.65
RM Score Ensemble 68.89 44.53 58.23 57.22
RM Agree.Ensemble 70.26 45.92 61.09 59.09
Online RM Ensemble 68.95 45.65 59.49 58.03
LASeR (Ours) 73.06 46.89 62.27 60.74
In instruction-following tasks, Fig. 10 shows that LASeR consistently outperforms all baselines, including Best RM (58.72%), Sequential RM Selection (63.72%), and Random RM Selection (70.61%). It achieves the highest win rate against RM Score Ensemble (73.27%) and surpasses Online RM Ensemble (65.41%), demonstrating its ability to generalize to different LLMs.
Figure 10: Length-controlled AlpacaEval win rates comparing LASeR against baselines on the instruction-following tasks on WildChat using Mistral-7B.
Table 11: LASeR outperforms baselines in long-context understanding tasks with Mistral-7B. Sequential RM selection is not applicable in this setting, as only inference is conducted. For QA and few-shot learning tasks, we report F1 scores, and for summarization, we report Rouge-L.
Method Single-Doc QA Multi-Doc QA Summarization Few-shot Learning
Base model 26.01 24.06 26.47 64.93
Best RM 28.93 27.93 30.42 68.34
Random RM Selection 27.44 25.38 27.19 66.72
Classifier RM Selection 27.59 25.69 27.97 67.04
RM Score Ensemble 26.75 25.71 28.17 66.97
RM Agree. Emsemble 27.96 26.60 28.27 67.23
Online RM Ensemble 27.55 25.42 28.63 66.81
LASeR (Ours) 29.14 27.80 30.08 68.47
In long-context understanding tasks, Table 2 highlights LASeR’s superior performance on long-context understanding tasks with Mistral-7B. LASeR achieves the highest scores for Single-Doc QA (29.14) and Multi-Doc QA (27.80), significantly outperforming the base model. In Few-Shot Learning, LASeR achieves 68.47, surpassing both the Best RM (68.34) and Online RM Ensemble (66.81). While its Rouge-L score for summarization (30.08) is slightly below Best RM (30.42), LASeR remains comparable. These results demonstrate the consistent performance of LASeR across the model family.
Generalization to the Base and RMs of Bigger Sizes.
To further analyze the generalization of LASeR to the base and RMs of bigger sizes, we add an experiment with Qwen2.5-32B [Qwen Team, 2024a] policy model and the following 4 similarly sized RMs from RewardBench: RISE-Judge-Qwen2.5-32B, Bagel-DPO-34B, Archangel-DPO-30B, Archangel-KTO-30B. We compare LASeR with the best RM and the RM Agreement Ensemble baselines, which are the second-best baselines for most of the benchmarks on Wildchat. Table 12 demonstrates that LASeR has higher win rates than baselines.
Table 12: Win rates of LASeR compared to Best RM and RM Agreement Ensemble baselines on Qwen-32B model.
Win Rates
LASeR vs. Best RM 54.44%
LASeR vs. RM Agree. Ensemble 55.78%
Generalization to the Number of RMs.
To study the generalization capability of LASeR across the number of RMs, we expand the candidate set of RMs with up to 4 more RMs from the RewardBench leaderboard, including Tulu-2-DPO-7B [Ivison et al., 2023], Zephyr-7B-Gemma [Tunstall and Schmid, 2024], Qwen1.5-MoE-A2.7B-Chat [Qwen Team, 2024b], Archangel-7B [Ethayarajh et al., 2024]. Fig. 11 shows that the accuracy remains consistent across all datasets as the number of RMs varies. StrategyQA remains near 85.9%, GSM8K around 74.8%, and MMLU close to 68.1%, with minimal fluctuations, indicating robust performance regardless of the number of RMs while RM Agreement Ensemble performance declines more steeply.
Generalization to Training Loss Functions.
In Sec. 3.1, we state that the choice of loss function used to train the LLM depends on the underlying task or domain. Nevertheless, we always use the training loss as the MAB reward to update the MAB’s parameters. Here we study the performance of LASeR and baselines with 4 different loss functions, NLL, DPO, NLL + DPO [Pang et al., 2024], and KTO [Ethayarajh et al., 2024], in the reasoning domain. Results in Table 13 show that training LLMs with multiple rewards using LASeR outperforms sequential RM selection by $2.4\%$, $1.7\%$, and $1.3\%$ when using NLL, DPO, NLL+DPO loss functions, respectively; while both methods yield comparable performance with KTO. Additionally, we found that the most effective training loss functions are NLL + DPO for StrategyQA, NLL for GSM8K, and KTO for MMLU. However, irrespective of the choice of the underlying loss function, LASeR is more effective at balancing and adaptively selecting from multiple RMs. Lastly, we compare LASeR with a variant in which we use $\mathrm{Acc}(y^{w})-\mathrm{Acc}(y^{l})$ as the MAB reward, which uses the ground-truth information about the final answer. We find that using the negative training loss of the LLM is more effective than using accuracy as the MAB reward.
@@ Appendix C Generalization Capabilities of LASeR
Table 13: Across different training loss functions, optimizing with multiple RMs via LASeR outperforms the sequential RM selection with Llama-3-8B. SQA denotes StrategyQA.
Loss Method SQA GSM8K MMLU Avg.
@@
LASeR w/. Acc 83.04 73.12 65.46 73.87
LASeR 85.96 74.75 68.24 76.24
Figure 11: LASeR’s performance is robust to adding weaker RMs to the candidate set to select from.
Generalization to OOD Tasks.
We first assess the generalization ability of our method by training models on specific datasets and evaluating their performance on out-of-distribution reasoning tasks. Specifically, we train the model on the StrategyQA and MMLU datasets and evaluate its generalization on the CommonsenseQA [CSQA; Talmor et al., 2019] dataset. Similarly, we train on GSM8K and test on MATH [Hendrycks et al., 2021c] to assess the model’s ability to generalize across different reasoning datasets. These tasks are designed to capture both general reasoning ability and OOD generalization across domains. We report the results in Table 14, where we find that across both Llama-3-8B and Mistral-7B models, models trained with LASeR yield the best average accuracy, beating training with the best RM by roughly $2\%$ (absolute) on CSQA with Llama-3-8B. On Mistral-7B, training with LASeR outperforms both training with single best RM and sequential RM selection by slightly over $1\%$.
@@
Method CSQA MATH Avg.
Llama-3-8B
SFT 65.64 29.13 47.39
Best RM 67.59 31.54 49.57
Avg. RM 67.16 30.36 48.76
Random RM Selection 68.31 30.21 49.26
Sequential RM Selection 67.73 30.25 48.99
LASER (Ours) 69.26 31.67 50.47
Mistral-7B
SFT 59.06 16.38 37.72
Best RM 60.46 18.08 39.27
Avg. RM 60.06 17.25 38.65
Random RM Selection 60.61 16.96 38.58
Sequential RM Selection 60.56 17.96 39.26
LASER (Ours) 61.65 18.97 40.31
Robustness to Noisy Rewards.
To examine the robustness of our method in the presence of noisy or irrelevant rewards, we conduct the following analysis using Llama-3-8B on GSM8K. We add varying amounts of Gaussian noise $\sigma$ to the rewards generated by RMs sampled from the distribution $\mathcal{N}(0,\sigma I)$, where $I$ is the identity matrix, to simulate noisy rewards when using RMs in out-of-distribution settings. In addition to LASeR using the LinUCB algorithm (c.f. Sec. 3.2), we also use Exp3 [Auer et al., 2002] designed for adversarial bandit settings. In Fig. 12, we find that even as the degree of noise in RM scores increases (from $\sigma\!=\!$ 0.1 to 0.4), LASeR’s selection strategy continues to perform robustly, mitigating the effects of noise compared to the sequential baseline. Specifically, LASeR has an average performance drop of only $0.55\%$ accuracy at a noise level of $\sigma\!=\!0.3$, whereas the sequential baseline suffers a $1.6\%$ accuracy drop (3 times as much) under the same conditions. Furthermore, LASeR using Exp3, the most noise-robust method, maintains consistent performance, with only a $0.26\%$ accuracy drop.
LASeR Training with Domain-specific Evaluation Metrics.
While recent works focus on building RMs that reflect preferences across domains, an extensive body of prior work develops a suite of evaluation metrics catered to specific domains such as reasoning [Golovneva et al., 2022, Prasad et al., 2023]. To show that LASeR can be used to select any kind of evaluation metric from a collection of metrics during training, in Table 15, we present results with training LLMs on model-based metrics from ROSCOE [Golovneva et al., 2022] by replacing RMs with informativeness, faithfulness, reasoning alignment, hallucination, common sense error, semantic, coherence and perplexity in Sec. 3. Llama-3-8B models trained using LASeR yield 1.62% accuracy improvement over baselines on average across three datasets. These results are also generalized to Mistral-7B, except for GSM8K, where we achieve comparable performance to the Base + Informativeness baselines. Note that the perplexity of most responses is nearly identical, making it difficult to distinguish between them, explaining why perplexity shows little to no improvement compared to the base model.
Method StrategyQA GSM8K MMLU Avg.
Llama-3-8B
SFT 80.41 69.43 65.66 71.83
Perplexity 80.55 69.21 65.62 71.79
Informativeness 82.87 73.55 66.69 74.37
Random RM Selection 82.72 70.93 66.10 73.25
Seq. RM Selection 83.15 73.38 66.17 74.23
LASER (Ours) 83.54 73.80 66.79 74.71
Mistral-7B
SFT 68.57 43.62 56.48 56.22
Perplexity 68.83 43.47 57.14 56.48
Informativeness 70.29 44.98 59.29 58.19
Random RM Selection 69.24 44.05 57.68 56.99
Seq. RM Selection 70.40 44.79 59.07 58.09
LASER (Ours) 70.91 44.93 59.63 58.49
Robustness to Underperforming Evaluation Metrics.
Similar to our analysis on noise in rewards in Sec. 5, we investigate how adding ROSCOE metrics with poor correlation to human-annotated labels in meta-evaluation by Golovneva et al. [2022] impacts the performance of Llama-3-8B on GSM8K. Once again, even with ROSCOE metrics, demonstrates LASeR can maintain consistent performance by adaptively prioritizing the most relevant reward signals, outperforming the sequential baseline, which fails to filter out irrelevant information effectively. Fig. 13 shows that as the number of irrelevant metrics increases, LASeR’s selection strategy continues to perform robustly. Specifically, LASeR has an average performance drop of only $0.13\%$, whereas the sequential baseline suffers a $2.15\%$ accuracy drop under the same conditions. Lastly, LASeR using Exp3 maintains a consistent performance level with a $0.4\%$ accuracy drop.
Appendix D Additional Discussion
LASeR with different “kinds” of RMs. In Sec. 4.2, we show that LASeR can choose from a set of candidate RMs, and our analysis in Fig. 12 highlights the fact that LASeR is robust to noisy RMs. In Appendix C, we show that LASeR can also be used with metric-based rewards [Golovneva et al., 2022]. These experiments reflect a conceptual split between the generator (the LLM) and the scorer (the RM or metric). Thus, LASeR is applicable to other settings that follow this paradigm, e.g., using an LLM-as-a-judge [Zheng et al., 2023], where LASeR could be used to choose between different judge models, prompts, or different combinations of RMs and metrics. However, consistent with the “self-preference” bias of LLMs [Panickssery et al., 2024], we caution that using an RM that is based on the same model as the LLM used for generating responses could lead to the MAB spuriously favoring certain RMs. We leave further study on extending LASeR to future work.
Quality of RMs used with LASeR. Methods that rely on RMs for scoring generally assume that these RMs have a strong correlation with human judgments. LASeR tempers this assumption in a number of ways: First, by ensembling multiple RMs, LASeR weakens the effect of noisy RMs; this can be seen in Fig. 12, where LASeR mitigates the negative impact of a noisy RM even as the level of noise is increased. Moreover, the fact that LASeR can select RMs at an instance level means that there need not be a single RM that always correlates well across all instances. However, LASeR does require at least one RM to be positively correlated with human judgments on each instance. If this assumption is not met (i.e., all RMs are poorly correlated across all instances), then optimizing for the RMs will yield poor results. Note that this holds true for any method optimizing for RMs. Because LASeR selects from multiple RMs, its contributions are complementary to developments in RMs, which can easily be integrated into LASeR, as well as improvements to preference optimization loss functions (see Appendix C). Such RM improvements are likely to be necessary as LLMs are deployed in domains that are out of scope for existing systems and domains with heterogeneous requirements (e.g., our generation domains in Sec. 4.2). In these cases, there will be no single “perfect” existing RM, and successful solutions will likely involve mixing multiple RMs. A core benefit of LASeR is its ability to automatically filter RMs; in Fig. 6 we see that utilization differs across domains. This allows users to avoid expensive experimentation with subsets of RMs: they can simply offload this task to LASeR, which will automatically select the more useful RM(s).
Appendix E Prompts
We provide the prompts used for our experiments as follows: